OpenStudy (anonymous):

Find the variation constant and an equation of variation when y varies directly as x and y=27 and x=9

OpenStudy (eyust707):

if y varies directly... y =kx replace with 27 and 9 27= k (9) thats your equation and k=3

OpenStudy (eyust707):

actually i think the equation would be: y = 3x

OpenStudy (anonymous):

So k=3 and y=3x?

OpenStudy (eyust707):

yep

OpenStudy (anonymous):

Thank you!

OpenStudy (eyust707):

got anymore?

OpenStudy (anonymous):

Oh, plenty. :D I'm doing a test. I'm doing the ones that I can do myself and putting the more complicated/ or ones i forgot how to do on here.

OpenStudy (eyust707):

ok let me know.. im hecka bored.. lol

OpenStudy (anonymous):

I'll attach one. Give me a sec.

OpenStudy (anonymous):

Zing.

OpenStudy (eyust707):

ok so we want (x-2)(x-2)=0 because if we set x = 2 its true

OpenStudy (eyust707):

just foil that out and there ya go =P

OpenStudy (anonymous):

k=y/x k=27/9=3

OpenStudy (anonymous):

Oh my, that was simple. lol I'm so tired I can't see straight. Here's another.

OpenStudy (anonymous):

5(6s+5t)^2

OpenStudy (eyust707):

notice with that one we can take out the t

OpenStudy (eyust707):

t (t^2-4t-45)

OpenStudy (eyust707):

then its just a normal one

OpenStudy (eyust707):

im thinkin -9 and +5...

OpenStudy (eyust707):

ill let you do the rest

OpenStudy (anonymous):

Awesome. Thank you!

OpenStudy (eyust707):

hehe no problem